TWiki home Point 1 Point 2 Point 3 Point 4 Point 5 Point 6 Point 7 Point 8 Minutes


Created by Main.MikeKoratzinos on 2007-06-18 - 18:15

Update on Powering tests sector 78 - Afternoon 18 June 2007

Progress during Monday 18.06.2007 from the 8:30 meeting till the 16:30 meeting

  • RQF.A78 and RQD.A78
    • Wait for green light from QPS - OK
    • PIC1 for both circuits: OK but several points to note:
      • PM problem noted: the data, although complete, does not display correctly in the PM viewer. Sandrine will look at data manually and Adriaan will correct the problem
      • after a discharge request from the PC, the PC fault and discharge request signals arrive out of sequence. According to Valerie this is understood (due to cable delays)
      • QPS_OK does not take into account the switches status allowing powering of the magnets with the switches open
    • Connection of the DC cables to the current leads has not happened yet. Will be done tomorrow at 7:30

  • RB.A78
    • QPS IST. All heaters charged (2 out of 700 needed intervention) Discharging started at 16:45 (will need 3-4hours to complete)

  • Matching Section
    • MPP presented the next tests to investigate the triple-quench event. It includes various ramping schemes, forced quench of Q4 and a boiloff test of DFBMA, to be completed during the next few days. Tonight: DFBMA boiloff and re-fill.

  • 600A correctors in DFBAO
    • ACR/MEL tests on the damaged cryo-instrumentation cards (RSF2.A78B2, RCS.A78B1) - ONGOING
    • If possible, solution of problem in RCO QPS communication -DONE
    • QPS Software upgrade - Will not done before tomorrow

  • Tests of the 60A CRYO_START/CRYO_MANTAIN signals - OK although for the 60A circuits all tests will have to be done in conjuction with the cryo operators with a manual CRYO_START/MAINTAIN

A.O.B.: Francois reminded the meeting that it is preferable that the cryo EIC is informed (as well as the cryo operators) of any pending tests.

Programme for tomorrow Tuesday 19 June 2007

  • RQF.A78 and RQD.A78
    • Connection of DC cables to the current leads
    • connection of a 60A converter to the RQD.A78 circuit from the UA83
    • PCC tests for RQD.A78 (MAXIMUM CURRENT: 350A) - Experts from AB/PO, PIC, QPS and Energy Extraction required
    • PIC2 tests for RQD.A78 in parallel with PCC for RQF.A78 (MAXIMUM CURRENT FOR BOTH TESTS: 350A)

  • RB.A78 QPS IST tests will continue during the whole day

  • Matching Section
    • Test 1 proposed by MPP (see slides attached)
    • Test 2 proposed by MPP, but:
      • If detection threshold can be increased up to 50 mV: full test
      • If detection threshold cannot be modified yet: full cycle without provoked quench at the end

  • 600A correctors in DFBAO:
    • ACR/MEL tests on the electronic cards will go on till RQD.A78 is released by PIC2
    • Software upgrade of the quench detectors

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-06-18 - 16:24

Beam Dump System - Installation & Commissioning - 05/06/2007

PDF RAT_20070605.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x06x18x16x21/RAT_20070605.pdf

Comments:

Edit, attach file or add comment


Created by Main.MariaPazCasasLino on 2007-06-18 - 12:34

Warm Magnets Commissioning at point 7. Minutes 18-06-2007

PDF RAT_20070618_WM7.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint7x2007x06x18x12x33/RAT_20070618_WM7.pdf

Comments:

Edit, attach file or add comment

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2007-06-01 - JacekSzkutnik
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    HCC All webs login